When an unbiased coin is tossed seven times, what is the probability of obtaining at least 4 heads?


1
Expert's answer
2021-05-07T09:52:44-0400

Let "X=" the number of heads: "X\sim Bin(n, p)"

Given "n=7, p=0.5"


"P(X\geq4)=P(X=4)+P(X=5)"

"+P(X=6)+P(X=7)"

"=\dbinom{7}{4}(0.5)^4(1-0.5)^{7-4}+\dbinom{7}{5}(0.5)^5(1-0.5)^{7-5}"


"+\dbinom{7}{6}(0.5)^6(1-0.5)^{7-6}+\dbinom{7}{7}(0.5)^7(1-0.5)^{7-7}"

"=\dfrac{1}{128}(35+21+7+1)=0.5"

The probability of obtaining at least 4 heads is 0.5.